OpenStudy (anonymous):

using L'hopital's rule evaluate lim x-->infinity (x-sqrt(x^2 +x)

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

do we have to use L`hopital rules ?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}(x-\sqrt{x^2+x})\] \[\lim_{x \rightarrow \infty}(x-\sqrt{x^2+x})*\frac{(x+\sqrt{x^2+x})}{(x+\sqrt{x^2+x})}\] \[\lim_{x \rightarrow \infty}\frac{-x}{(x+\sqrt{x^2+x})}\] \[\lim_{x \rightarrow \infty}\frac{-x}{x(1+\sqrt{1+\frac1{x^2}})}\] \[\lim_{x \rightarrow \infty}\frac{-1}{(1+\sqrt{1+\frac1{x^2}})}=\frac{-1}2\]